# Separating the wave and particle attributes of two entangled photons

### v2

@@ III Separation of wave-particle properties of two entangled photons
Next, by adjusting the beam-splitters $BS_{3}$ and $BS_{4}$, respectively, we make the states $|R_{1}L_{2}\rangle$ and $|L_{1}R_{2}\rangle$ go through $X_{3}$ and $X_{4}$, such that the detectors $D_{1}$ and $D_{2}$ will not click. The $X_{3}=X_{4}=|W\rangle\langle W|$ operators transmit only the wave state, and this action guarantees that detectors $D_{3}$ and $D_{4}$ will not click. Finally, by adjusting the beam splitter $BS_{5}$, the state $|R_{1}L_{2}\rangle+|L_{1}R_{2}\rangle$ reaches the detector $D_{5}$. Overall, we conclude that if and only if the detector $D_{5}$ clicks we have made a successful post-selection. Otherwise, if any other detector clicks post-selection fails, and we fail to separate the wave and particle attributes of the two entangled photons.
The operators that measure whether the wave and particle attributes of two entangled photons are present in the left and right arms is defined as
@@
with $\nu\in\{W,W^{\prime},P,P^{\prime}\}$, $\mu_{i}\in\{L_{i},R_{i}\}$ and $i\in\{1,2\}$. After preparing the pre-selected state, we choose to post-select by $|\psi_{f}\rangle$ and perform weak measurement for observables $\Pi_{\nu}^{\mu_{i}}$. By substituting the above-defined pre- and post-selected states into Eq. (3), we find the corresponding weak values of the observables $\Pi_{\nu}^{\mu_{i}}$, i.e.,
$\langle\Pi_{W}^{\mu_{i}}\rangle$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{R_{i}\mu_{i}},$ (23a)
$\langle\Pi_{P}^{\mu_{i}}\rangle$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{L_{i}\mu_{i}},$ (23b)
$\langle\Pi_{W^{\prime}}^{\mu_{i}}\rangle$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{L_{i}\mu_{i}},$ (23c)
$\langle\Pi_{P^{\prime}}^{\mu_{i}}\rangle$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{R_{i}\mu_{i}}.$ (23d)
@@
If $\alpha\neq 0,\frac{\pi}{2}$, these weak values indicated that the wave and particle attributes of two entangled photons are perfectly separated, and in each one of the four arms we only have one attribute (wave or particle ) of the two photons. In the particular case, $\alpha=\frac{\pi}{4},$ the weak values for the wave and particle attributes may be written as $\langle\Pi_{P}^{L_{1}}\rangle=\langle\Pi_{W}^{R_{1}}\rangle=\frac{1}{2}$, $\langle\Pi_{P^{\prime}}^{R_{2}}\rangle=\langle\Pi_{W^{\prime}}^{L_{2}}\rangle=\frac{1}{2}$. This means that for $\alpha=\frac{\pi}{4}$ case, half of the particle attributes of the two entangled photons occurrs in $L_{1}$ and $R_{2}$ arms and half of the wave attributes are present in $R_{1}$ and $L_{2}$ arms of the interferometer, respectively. Furthermore, it is clear that these weak values cannot take arbitrary large values, and always obey the Bohr’s complementarity principle, i.e.,
$\langle\Pi_{P}^{L_{1}}\rangle+\langle\Pi_{W}^{R_{1}}\rangle$ $=1,$ (24)
$\langle\Pi_{P^{\prime}}^{R_{2}}\rangle+\langle\Pi_{W^{\prime}}^{L_{2}}\rangle$ $=1.$ (25)
IV conclusion
In this paper, we have investigated the possibility of spatially separating the wave and particle features of two entangled photons. Our work is based on measuring the weak values of two-level quantum systems by selecting appropriate pre- and post-selection states in order to generate polarization-path like quantum Cheshire cats.
@@
We found that properly pre- and post-selecting the two-photon state, the wave and particle attributes of two entangled photons can be spatially separated, while keeping the validity of Bohr’s complementarity principle. We also proposed a feasible setup for the implementation of our thought experiment on a quantum optical platform.
@@
The quantum entanglement is a cornerstone of quantum informatics and it has an irreplaceable role in quantum precision measurements. In particular, in quantum metrology one would generally require that the measurement of a physical quantity on an entangled state does non disturb to other variable. Our results may foster the exploration of schemes to separate entangled observables and enhance precision of quantum metrological protocols. At the same time, in the emerging field of counterfactual communication (Salih et al., 2013; Li et al., 2015), information may be exchanged even without the transmission of quantum state. In this framework, entanglement also plays a vital role, but the basic principles of communication is still unclear. Also in this case, our proposal may promote the understanding of the fundamental mechanism behind counterfactual communication.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 12365005).

### v4

@@ III Separation of wave-particle properties of two entangled photons
Next, by adjusting the beam-splitters $BS_{3}$ and $BS_{4}$, respectively, such that states $|R_{1}\rangle$ and $|L_{1}\rangle$ ($|L_{2}\rangle$ and $|R_{2}\rangle$) totally transmitted (reflected) to guarantee $|R_{1}L_{2}\rangle$ and $|L_{1}R_{2}\rangle$ go through $X_{3}$ and $X_{4}$. Under this action the detectors $D_{1}$ and $D_{2}$ both will not click. In this step, the modified form of $|\psi_{f^{\prime}}\rangle$ can be written as
$|\psi_{f^{\prime}}\rangle\overset{BS_{3},BS_{4}}{\longrightarrow}|\psi_{f^{\prime\prime}}\rangle=\frac{1}{\sqrt{2}}\left(|R\rangle+|L\rangle\right)|\mathcal{\boldsymbol{W}}\rangle.$ (21)
Here, we denote $|R_{1}L_{2}\rangle\overset{BS_{3}}{\longrightarrow}|R\rangle$, $|L_{1}R_{2}\rangle\overset{BS_{4}}{\longrightarrow}|L\rangle$ and $|\mathcal{\boldsymbol{W}}\rangle=|WW^{\prime}\rangle$. The $X_{3}=X_{4}=|W\rangle\langle W|$ operators transmit only the wave state, and this action guarantees that detectors $D_{3}$ and $D_{4}$ will not click. The beam splitter $BS_{5}$ is chosen as $|R\rangle\overset{BS_{5}}{\longrightarrow}\frac{1}{\sqrt{2}}\left(|R\rangle+|L\rangle\right)$, $|L\rangle\overset{BS_{5}}{\longrightarrow}\frac{1}{\sqrt{2}}\left(|R\rangle-|L\rangle\right)$, such that
$|\psi_{f^{\prime\prime}}\rangle\overset{BS_{5}}{\longrightarrow}|\Psi_{f}\rangle=|R\rangle|\mathcal{\boldsymbol{W}}\rangle.$ (22)
Hence, we can confirm that detector $D_{5}$ is click with certainty. Contrarily, if and only if the detector $D_{5}$ click with $100\%$ probability, then our schematics also can prove that the postslected state should be chosen to $|\psi_{f}\rangle$.
After choosing the pre- and post-selected state of our scheme, we have to extract the weak values to achieve our goal. The operators that measure whether the wave and particle attributes of two entangled photons are present in the left and right arms is defined as
@@
with $\nu\in\{W,W^{\prime},P,P^{\prime}\}$, $\mu_{i}\in\{L_{i},R_{i}\}$ with $i=1,2$. Above every operators has its meaning, such as $\Pi_{W}^{L_{1}}=|L_{1}\rangle\langle L_{1}|\otimes|W\rangle\langle W|$ can describe the findings of the wave attribute of first photon on the left arm and $\Pi_{P^{\prime}}^{R_{2}}=|R_{2}\rangle\langle R_{2}|\otimes|P^{\prime}\rangle\langle P^{\prime}|$ can describe the findings of the particle attribute of second photon on the right arm, etc.
@@
By substituting the above-defined pre- and post-selected states $|\psi_{f}\rangle$ and $|\psi_{f}\rangle$ into Eq. (3), we obtain the corresponding weak values of the observables $\Pi_{\nu}^{\mu_{i}}$, i.e.,
$\langle\Pi_{W}^{\mu_{i}}\rangle_{w}$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{R_{i}\mu_{i}},$ (24a)
$\langle\Pi_{P}^{\mu_{i}}\rangle_{w}$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{1i}\delta_{L_{i}\mu_{i}},$ (24b)
$\langle\Pi_{W^{\prime}}^{\mu_{i}}\rangle_{w}$ $=\frac{\cos\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{L_{i}\mu_{i}},$ (24c)
$\langle\Pi_{P^{\prime}}^{\mu_{i}}\rangle_{w}$ $=\frac{\sin\alpha}{\cos\alpha+\sin\alpha}\delta_{2i}\delta_{R_{i}\mu_{i}}.$ (24d)
@@
This is the main result of our study, and they can characterize the exact values of operators $\Pi_{\nu}^{\mu_{i}}$ in the region from $|\psi_{i}\rangle$ to $|\psi_{f}\rangle$. If $\alpha\neq 0,\frac{\pi}{2}$, these weak values indicated that the wave and particle attributes of two entangled photons are spatially separated, and in each one of the four arms we only have one attribute (wave or particle ) of the two photons. In the particular case, $\alpha=\frac{\pi}{4},$ the weak values for the wave and particle attributes may be written as $\langle\Pi_{P}^{L_{1}}\rangle_{w}=\langle\Pi_{W}^{R_{1}}\rangle_{w}=\frac{1}{2}$, $\langle\Pi_{P^{\prime}}^{R_{2}}\rangle_{w}=\langle\Pi_{W^{\prime}}^{L_{2}}\rangle_{w}=\frac{1}{2}$. This means that for $\alpha=\frac{\pi}{4}$ case, half of the particle attributes of the two entangled photons occurred in $L_{1}$ and $R_{2}$ arms and half of the wave attributes are present in $R_{1}$ and $L_{2}$ arms of the interferometer, respectively. Furthermore, it is clear that these weak values cannot take arbitrary large values, and always obey the Bohr’s complementarity principle, i.e.,
$\langle\Pi_{P}^{L_{1}}\rangle_{w}+\langle\Pi_{W}^{R_{1}}\rangle_{w}$ $=1,$ (25)
$\langle\Pi_{P^{\prime}}^{R_{2}}\rangle_{w}+\langle\Pi_{W^{\prime}}^{L_{2}}\rangle_{w}$ $=1.$ (26)
In this proposal, the correctly choosing of post-selected state is very crucial and we only focus on the cases of $D_{5}$ click with certainty. Otherwise, if any other detector clicks, the post-selection fails, and we fail to separate the wave and particle attributes of two entangled photons.
IV CONCLUSION AND DISCUSSION
@@
In this paper, we have investigated the possibility of spatially separating the wave and particle features of two entangled photons. Our work is based on measuring the weak values of system by selecting appropriate pre- and post-selection states in order to generate polarization-path quantum Cheshire cat states of two entangled photons. We found that by properly choosing the pre- and post-select states of system, the wave and particle attributes of two entangled photons can be spatially separated, while keeping the validity of Bohr’s complementarity principle. We also proposed a feasible setup for the implementation of our thought experiment with quantum optical platform.
@@
As indicated in our result, we can control wave-particle attributes of two entangled photons by adjusting the optical axis angle $\alpha$ of HWP. If $\alpha=0$ or $\alpha=\frac{\pi}{2}$, then we only have pure wave attributes or pure particle attributes of two photons and separate them spatially. Thus, it is very curious issue to do double- slit experiment by using pure particle attributes of entangled photons (Hong and Noh, 1998). Furthermore, in the emerging field of counterfactual communication (Salih et al., 2013; Li et al., 2015), information may be exchanged even without the transmission of quantum systems. In this framework, entanglement also plays a vital role, but the basic principles of such communication is still unclear. Thus, it may interesting to use pure wave attributes of entangled photons for understanding of the fundamental mechanism behind counterfactual communication. At the same time, our results may also provide an alternative method to enhance the precision of quantum metrological protocols. In particular, in quantum metrology one would generally require that the measurement of a physical quantity on an entangled state does non disturb to other variable. Thus, the separation of entangled observables by QCC and implement the related precision measurement may dramatically increase its efficiency.
Another interesting point of our scheme is that it can scale up to more multi-entangled photons cases. Since the parallel assembly of $N$ single-photon WP toolboxes allows us to generate $N$-photon wave-particle entangled states (Rab et al., 2017), we may realize the spatial separation of wave and particle attributes of $N$-photon wave-particle entangled states by expand our schematics.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 12365005)